OpenStudy (anonymous):

find the area of a triangle when A=22 degrees B=105 degrees and b=14

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@Jeremy0203

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

Hi, Area = \[\frac{ 1 }{ 2 }b * h\]b = base h = vertical height A=22 degrees B=105 degrees and b=14. However, you will want to use Heron's Formula for this, since you are given 3 angles, and not sides. You can calculate the area of a triangle if you know the lengths of all three sides, using a formula that has been know for nearly 2000 years. It is called "Heron's Formula" after Hero of Alexandria (see below) Just use this two step process: Calculate "s" (half of the triangles perimeter) using:\[s = \frac{ a + b + c }{ 2 }\] Then calculate the Area using: \[A = \sqrt{s(s-a)(s-b)(s-c)}\] Example: What is the area of a triangle where every side is 5 long? Step 1: s = (5+5+5)/2 = 7.5 Step 2: A = √(7.5 × 2.5 × 2.5 × 2.5) = √(117.1875) = 10.825... ______________________________ Drawing back: A=22 degrees B=105 degrees and b=14 If you plug that in to the formula I gave you, you'll see A = ???

OpenStudy (anonymous):

30.4?

OpenStudy (anonymous):

@Compassionate